Back to main Archive News page.
Fri 31 Mar 2006 - Central servers at risk, otherwise all systems running without major problems
Several central systems will be at risk this weekend due to urgent electrical work. The affected systems include the central web server www.lancs.ac.uk, student mail systems (Wing, and mail.lancs.ac.uk) and the central unix service. The systems will be running on a temporary power supply, from 16:00 on Friday until 10:00 on Monday.
Thu 30 Mar 2006 - All systems ran without major problems
Wed 29 Mar 2006 - Exchange Email problem resolved, Central File Server maintenance, and Virus Attack
Yesterday's problems with the Exchange Server were resolved at 08:30. The Email team do not believe that any email or other Outlook items were lost as a result of the outage.
The central file servers were briefly unavailable for the application of security updates between 13:00 and 18:00.
On 29th March, several PCs were attacked by an unknown virus. As a result of the attack, a number of PCs had to be rebuilt causing inconvenience to all concerned. The infected PCs used 'shared folders' to share data. It is suspected that the virus spread through 'shared folders'. ISS strongly recommends that users use the 'Departmental Filestore' system to share data.
Tue 28 Mar 2006 - Exchange Email was Unavailable
Due to hardware problems, the Exchange email system was unavailable. This affected all users of Outlook (including Outlook Web Access and Outlook Express). ISS technical staff worked hard throughout the day to restore the service.
UPDATE 15:30: The EXCHANGE-BE2 server is now available and users whose mailboxes are held on it can work normally (approximately half the users, plus Public Folders). We estimate that service on the EXCHANGE-BE1 server will be resumed either late this evening or early tomorrow morning. The delay is due to a restore from tape and extensive system checking being necessary.
UPDATE 08:30 Wednesday 29th March: The EXCHANGE-BE1 server is back in service, now making the Exchange email system fully operational. The system will be slow for some time after this as the backlog of messages is cleared. The email team believe that no email or other Outlook items were lost as a result of the outage.
Tue 28 Mar 2006 - Network Renumbering & Electrical work
07:00-09:00: Network renumbering was carried out in InfoLab, and various equipment was moved to new electrical supplies.
Mon 27 Mar 2006 - All systems ran without problems
Fri 24 Mar 2006 - All systems ran without problems
Thu 23 Mar 2006 - WiNet (Wireless Network) problems
An issue with the vendor software on the Wireless gateway (BlueSocket), caused problems throughout the day. The software has now been updated and the facility is fully serviceable once again.
Wed 22 Mar 2006 - All systems ran without problems
Tue 21 Mar 2006 - Network maintenance, otherwise all systems ran without major problems
07:00-09:00 During the 'at risk' period, further work to improve performance in Engineering was performed. The work replaced a core router which is believed to be causing some performance problems seen last week.
07:00-09:00 Due to problems seen last week during routine work, the suppliers of our core network recommended we upgrade the software version we are using. This has now been done.
Mon 20 Mar 2006 - All systems ran without problems
Fri 17 Mar 2006 - Internet problems
Due to a failure of JANET Internet circuits at Warrington, we were without
Internet access. This also affected web-based print credit payments. Worldcom
Engineers are working on the fibre circuits to locate the problem.
UPDATE: Sat 07:45: Unfortunately the problem within the JANET network does not appear to have been resolved overnight, we're currently awaiting an update from the JANET on-call engineers.
UPDATE: Sat 07:55: One of the 3 affected JANET circuits has now been restored, bringing back our internet connectivity. Apologies for any inconvenience.
UPDATE: Sat 15:29: All 3 of the JANET circuits have now been repaired. As a result normal service and performance to the internet should have been restored.
Thu 16 Mar 2006 - All systems running without major problems
Wed 15 Mar 2006 - All systems ran without major problems
Tue 14 Mar 2006 - All systems ran without major problems
08:00-10:00 During the 'at risk' period, work to resolve performance problems in the Engineering building and resilience to Fylde was undertaken.
Mon 13 Mar 2006 - All systems ran without major problems
Fri 10 Mar 2006 - All systems ran without problems
Thu 9 Mar 2006 - All systems ran without problems
Wed 8 Mar 2006 - Exchange Server Maintenance, otherwise all systems ran without problems
The Exchange Email system was unavailable for a brief interval between 3PM and 3:15PM to correct a system error.
Tue 7 Mar 2006 - All systems ran without problems
Mon 6 Mar 2006 - All systems ran without problems
Fri 3 Mar 2006 - All systems ran without problems
Thu 2 Mar 2006 - All systems ran without problems. County College Lab closed today
The ISS PC Lab at County College was closed on Thursday 2nd March for essential cabling work.